# ICONS: Influence Consensus for Vision-Language Data Selection

### v2

@@ Abstract
Visual Instruction Tuning typically requires a large amount of vision-language training data. This data often containing redundant information that increases computational costs without proportional performance gains. In this work, we introduce ICONS, a gradient-driven Influence CONsensus approach for vision-language data Selection that selects a compact training dataset for efficient multi-task training. The key element of our approach is cross-task influence consensus, which uses majority voting across task-specific influence matrices to identify samples that are consistently valuable across multiple tasks, allowing us to effectively prioritize data that optimizes for overall performance. Experiments show that models trained on our selected data (20% of LLaVA-665K) achieve 98.6% of the relative performance obtained using the full dataset. Additionally, we release this subset, LLaVA-ICONS-133K, a compact yet highly informative subset of LLaVA-665K visual instruction tuning data, preserving high impact training data for efficient vision-language model development.
Figure 1: Influence Consensus for Vision-Language Data Selection. (Left) Given a large scale visual instruction tuning dataset (LLaVA-665K [23]), our method computes an influence score of each training example on each target task, aiming to select examples with high influence across multiple tasks. Our selected compact 20% subset (LLaVA-ICONS-133K) contains data points that achieve influence consensus across tasks, as indicated by the green check marks showing high influence ranking on corresponding tasks. (Right) The radar plot shows the relative performance on various benchmarks comparing training with LLaVA-665K (the full dataset) and LLaVA-ICONS-133K (our selected subset), where the selected subset achieves comparable performance to the full dataset.
@@ 1 Introduction
Visual instruction tuning has become a critical stage in developing Multimodal LLMs [23, 22], enabling these models to effectively interpret and respond to language instructions based on visual content. Current approaches typically rely on large instruction tuning datasets (e.g., LLaVA 1.5 [22] and Cambrian [39], which use 665K and 7M instruction tuning data points, respectively). These large datasets, while effective, pose practical challenges for research iteration and model development: extended development cycles due to lengthy training times [3, 15], increased storage and memory demands for managing large-scale datasets [36, 8], and substantial computational resources required for training [27, 40]. This motivates a fundamental question:
@@
Can we identify a compact subset of training data that preserves model capabilities while enabling faster experimentation?
Prior work has explored various approaches to data selection, including gradient-based methods [42, 6], influence functions [43, 17], and diversity sampling [44, 4]. These methods typically either focus on optimizing for single tasks in isolation, or maximizing source data diversity without considering target task distributions. However, selecting data for visual instruction tuning requires careful consideration of performance across diverse downstream tasks. Simply selecting the most influential samples for an individual task risks creating a dataset that performs well on that specific benchmark but fails to develop broader capabilities. Rather than focusing on task-specific performance, we aim to identify universally valuable training samples by aggregating gradient-based influence measures across diverse tasks using a voting mechanism, effectively selecting data that contributes to broad model capabilities.
@@
We introduce ICONS (Influence CONsensus vision-language data Selection), a method that builds upon the gradient-based influence estimation approach from LESS [42]. Given access to validation data for each target task, our method: (1) computes first-order gradient influence scores to measure how each training sample impacts task-specific performance, and (2) uses influence consensus through majority voting to identify training samples that show consistent positive value across multiple tasks. This consensus-based mechanism identifies universally valuable training examples: while some samples might be highly influential for individual tasks, we prioritize those that demonstrate broad utility across the task spectrum. While the computational cost of influence estimation is expensive, this front-loaded, one-time investment yields a standardized, compact dataset that can significantly accelerate iteration and development of both multimodal data and models.
@@
Using ICONS, we create LLaVA-ICONS-133K, an automatically curated 20% subset of the LLaVA dataset [23]. This compact dataset maintains 98.6% of the original performance across multiple vision-language tasks, providing a 2.8 percentage point improvement over randomly selecting the same-sized subset (95.8%) and eliminating two-thirds of the performance drop from shrinking the training data. Moreover, our ICONS outperforms all baselines across different selection ratios, and remarkably achieves above-full-dataset performance, surpassing the original dataset at a 60% selection ratio. Importantly, LLaVA-ICONS-133K shows strong transferability, maintaining 95.5-113.9% relative performance across diverse unseen tasks such as chart understanding and infographic comprehension, suggesting that ICONS identifies fundamentally valuable training examples rather than task-specific patterns. We summarize our key contributions:
@@
    We propose ICONS, a simple yet effective approach for multi-task vision-language data selection that identifies broadly valuable training samples through majority voting across task-specific gradient influence scores.
@@
    Our voting-based aggregation consistently outperforms all baselines (§3.2) and, most notably, exceeds 102% of the full dataset performance at 60% selection ratio (§3.3). We further analyze different influence aggregation approaches and show the effectiveness of our consensus-aware selection over direct-merge alternatives (§3.5).
@@
    We release LLaVA-ICONS-133K, a compact 20% subset of LLaVA-665K, achieving near-full performance (98.6%), transferring well to unseen tasks (§3.4), and serving as a standardized training set for resource-efficient development of multimodal models.
@@ 2 Influence Consensus for Vision-Language Data Selection
We propose a consensus-driven, gradient-based data selection framework for visual instruction tuning datasets. We formalize the problem setup in §2.1 and establish gradient-based influence estimation preliminaries in §2.2. Our two-stage data selection framework consists of: first, the specialist stage (§2.3), which computes task-specific influence scores, followed by the generalist stage (§2.4), which builds cross-task consensus through voting-based aggregation.
@@ 2.1 Problem Formulation
Given a large-scale visual instruction tuning dataset $\mathcal{D}=\{({\bm{x}}_{i},{\bm{I}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ containing $N$ samples, where each data point ${\bm{z}}_{i}=({\bm{x}}_{i},{\bm{I}}_{i},{\bm{y}}_{i})$ includes natural language instruction ${\bm{x}}_{i}$ and an image ${\bm{I}}_{i}$, with corresponding target response ${\bm{y}}_{i}$ [^1] The framework supports multi-turn conversational data, yet we formalize the problem setup for single-turn instruction-tuning for clarity and simplicity., and given access to validation data $\mathcal{V}_{k}$ for each downstream task $T_{k}\in\mathcal{T}=\{T_{1},...,T_{K}\}$, our goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes model performance across multiple downstream tasks:
$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=m}\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k}),$ (1)
$\text{Rel}(f_{\mathcal{S}},T_{k})=\frac{\text{Score}(f_{\mathcal{S}},T_{k})}{\text{Score}(f_{\mathcal{D}},T_{k})},$ (2)
where $f_{\mathcal{S}}$ and $f_{\mathcal{D}}$ denote models trained on subset $\mathcal{S}$ and full dataset $\mathcal{D}$, respectively. $\text{Score}(f,T_{k})$ is the task-specific evaluation score achieved by model $f$ on task $T_{k}$. We define the average relative performance across all tasks as Rel. = $\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k})/K$. Rel. quantifies the subset-trained model’s performance relative to that of the model trained on the entire dataset, with values close to 1 indicating that the subset maintains the performance of full training [18]. Our objective is to find a subset where Rel. $\approx 1$, i.e., the model trained on the selected subset achieves performance comparable to using the full dataset across all tasks.
@@ 2.2 Preliminaries
Building on our problem formulation in §2.1, we formalize how to estimate the influence of training samples on downstream task performance. Specifically, since our goal is to maximize $\text{Rel}(f_{\mathcal{S}},T_{k})$ across tasks as defined in Eqn. 2, we need an efficient way to estimate how each training sample contributes to the $\text{Score}(f_{\mathcal{S}},T_{k})$ term in the numerator. Following [34, 42], we estimate how individual training samples $z_{i}\in\mathcal{D}$ affect validation loss by measuring their gradient alignment with reducing validation loss on $\mathcal{V}_{k}$, which directly impacts task-specific evaluation scores. Denote a training data point as ${\bm{z}}$ and a validation data point as ${\bm{z}}^{\prime}$ from validation set $\mathcal{V}_{k}$ for task $T_{k}$ When training with SGD and batch size 1, using data point ${\bm{z}}$ at timestep $t$ leads to a model update ${\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}\nabla\ell({\bm{z}};{\bm{\theta}}_{t})$, where $\eta_{t}$ is the learning rate. To reduce the computational cost, we use the first-order Taylor expansion to estimate the loss on a given validation data point ${\bm{z}}^{\prime}$ at time step $t+1$ as:
@@
$\mathcal{I}_{t}({\bm{z}}\rightarrow{\bm{z}}^{\prime})$ $=\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t+1})-\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t})$
$\approx-\eta_{t}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
which we refer to as an influence score. For our case of visual instruction tuning [23] which uses single-epoch training, the influence score is:
$\mathcal{I}_{\text{single}}({\bm{z}}\rightarrow{\bm{z}}^{\prime})=\eta\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
where $\eta$ is the learning rate used during training. This can be generalized to multi-epoch training by accumulating the influence scores over all $n$ epochs:
$\mathcal{I}_{\text{total}}({\bm{z}}\rightarrow{\bm{z}}^{\prime})=\sum_{i=1}^{n}\eta_{i}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{i}),\nabla\ell({\bm{z}};{\bm{\theta}}_{i})\rangle,$
where $\eta_{i}$ is the learning rate used during the $i$th epoch. The gradient-based selection approach selects training samples $\{{\bm{z}}\}$ that maximize the gradient inner product $\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle$ [^2] In practice, we use cosine similarity instead of direct inner products to avoid biasing selection toward shorter sequences, since gradient norms tend to be inversely correlated with sequence length as noted in [42]. through a greedy, first-order approximation, which leads to larger reductions in validation loss for point ${\bm{z}}^{\prime}$. While this approach omits second-order terms compared to influence functions [17], it provides an efficient approximation for ranking the impact of training samples [11, 42].
@@ 2.3 Specialist: Individual Task Influence Ranking
To rank the influence of training data for each target task, we compute the influence score of each training data point on a validation set that represents the target task distribution. Following LESS [42], the process involves three steps: (1) training the model on 5% randomly selected data as a lightweight warm-up phase to initialize visual instruction-following capabilities, (2) computing gradients for training and validation data and compressing the gradients via random projection, and (3) computing the influence score to quantify the impact of each training data on validation set.
Step 1: Warm-up Training.  Following LESS [42], we first perform a brief warmup phase using Low-Rank Adaptation (LoRA) [12] on a small random subset $\mathcal{D}_{\mathrm{warmup}}\subset\mathcal{D}$ (5% of the training data): $f_{\text{warmup}}=\text{LoRA}(f_{\text{base}},\mathcal{D}_{\text{warmup}})$. This warmup phase allows the model to develop basic instruction-following capabilities.
@@
where $y_{i}$ and $y^{\prime}_{j}$ are the targets for $z_{i}$ and $z^{\prime}_{j}$, respectively. In order to reduce computational and storage overhead, we apply random projection to the gradient feature: $\tilde{g}_{i}=Rg_{i}$ and $\tilde{g}^{\prime}_{j}=R^{\prime}g^{\prime}_{j}$, where $R,R^{\prime}\in\mathbb{R}^{d^{\prime}\times d}$ is a random projection matrix with $d^{\prime}\ll d$ that preserves inner products with high probability [14]. We further normalize the projected gradient features, $\tilde{g}_{i}=\frac{\tilde{g}_{i}}{\|\tilde{g}_{i}\|_{2}},\tilde{g}^{\prime}_{j}=\frac{\tilde{g}^{\prime}_{j}}{\|\tilde{g}^{\prime}_{j}\|_{2}}$ to prevent bias from sequence length differences [42].
Step 3: Influence Matrix Computation.  We compute the influence matrix $I\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}^{k}|}$ where each entry $I_{ij}=\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$ is the influence of training data ${\bm{z}}_{i}$ on validation data ${\bm{z}}^{\prime}_{j}$, and the average influence of training data ${\bm{z}}_{i}$ on the target task $k$ is then computed as:
@@
This influence estimation process provides a task-specific ranking for the training set $\mathcal{D}$ with respect to task $\mathcal{T}_{k}$, where a higher influence score $\bar{I}_{i}$ suggests a higher influence for the target task. However, recall that our ultimate goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes $\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k})/K$ across all tasks. We address this disconnection between task-specific rankings and overall dataset optimization by proposing a voting-based generalist approach; this approach aggregates influence scores for each target task to identify the most broadly impactful training data.

### v3

@@ Abstract
Training vision-language models via instruction tuning often relies on large mixtures of data spanning diverse tasks and domains. However, these mixtures frequently include redundant information, increasing computational costs without proportional performance gains—necessitating more effective data selection strategies. Existing methods typically rely on task-agnostic heuristics to estimate data importance or focus on optimizing single tasks in isolation, limiting their effectiveness in multitask settings. In this work, we introduce ICONS, a gradient-based Influence CONsensus approach for vision-language data Selection. Our method leverages first-order training dynamics to estimate the influence of individual training examples on validation performance and aggregates these estimates across tasks via majority voting over task-specific influences. This cross-task consensus identifies data points that are consistently valuable across tasks, enabling us to prioritize examples that drive overall performance. The voting-based design further mitigates issues such as score calibration and outlier sensitivity, resulting in robust and scalable data selection for diverse multitask mixtures. With only 20% of the data from LLaVA-665K and Cambrian-7M, our selected subsets retain 98.6% and 98.8% of the performance achieved with full datasets—and can even surpass full-data training at a 60% selection ratio on LLaVA-665K. Our approach also generalizes to unseen tasks and architectures, demonstrating strong transfer. We release two compact, high-utility subsets—LLaVA-ICONS-133K and Cambrian-ICONS-1.4M —preserving impactful training examples for efficient and scalable vision-language model development.
@@ 1 Introduction
Visual instruction tuning is a crucial step in training multimodal language models liu2024improved ; llava , enabling them to follow language instructions grounded in visual content. Recent approaches rely on large-scale datasets such as LLaVA-665K liu2024improved and Cambrian-7M tong2024cambrian , which contain 665K and 7M examples, respectively. While effective, these datasets introduce significant barriers to iteration and deployment: prolonged training times brown2020language ; kaplan2020scaling , high storage demands rajbhandari2020zero ; gadre2024datacomp , and substantial compute costs maini2024rephrasing ; wu2022sustainable . Moreover, not all examples contribute equally to all tasks—naively scaling up diverse data mixtures can introduce redundancy and inefficiency. This raises a fundamental question:
{mdframed}
[leftmargin=0pt, rightmargin=0pt, innerleftmargin=10pt, innerrightmargin=10pt, skipbelow=0pt]
@@
Can we identify a compact, multitask-effective subset of training data that preserves model capabilities arcoss tasks while enabling faster experimentation?
Prior work has explored various data selection strategies, including gradient-based approaches xia2024less ; deng2024influential , influence functions yang2022dataset ; koh2017understanding , and diversity-based sampling yu2024diversify ; bukharin2023data . However, many of these methods either optimize for single tasks in isolation or maximize source diversity without aligning to downstream needs. In multitask visual instruction tuning, this is particularly limiting: optimizing for one task may hurt generalization, and task-agnostic diversity may dilute impact. Rather than selecting
Figure 1: Influence consensus for vision-language data selection. (Left) Given a large scale visual instruction tuning dataset (LLaVA-665K), our method uses majority voting across task-specific influence scores to identify training samples that are consistently influential across multiple tasks, forming a compact 20% subset (LLaVA-ICONS-133K) with data points achieving influence consensus. (Right) The radar plot compares performance between LLaVA-665K and our selected subset, showing the selected subset achieves comparable results to the full dataset.
data based on per-task influence, we aim to identify samples that are broadly useful—training examples that consistently contribute across tasks. To do this, we aggregate gradient-based influence scores using a simple yet effective majority voting scheme.
@@
We introduce ICONS (Influence CONsensus vision-language data Selection), a method that builds upon the gradient-based selection approach LESS xia2024less . Given access to validation data for each target task, our method: (1) computes first-order gradient influence scores to measure how each training sample impacts task-specific performance, and (2) uses influence consensus through majority voting to identify training samples that show consistent positive value across multiple tasks. This consensus-based mechanism identifies universally valuable training examples: while some samples might be highly influential for individual tasks, we prioritize those that demonstrate broad utility across the task spectrum. While the computational cost of influence estimation is expensive, this front-loaded, one-time investment yields a standardized, compact dataset that can significantly accelerate development of multimodal models, and enables reusable gradient datastores that amortize costs across iterations and deliver long-term savings.
@@
Using ICONS, we create LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, automatically curated 20% subsets of the LLaVA-665K dataset liu2024improved and Cambrian-7M tong2024cambrian dataset, respectively. These compact datasets maintain 98.6% and 98.8% of their original performance across multiple vision-language tasks, providing significant improvements over randomly selecting same-sized subsets (95.8% and 95.4%) and eliminating approximately two-thirds of the performance drop from shrinking the training data. Moreover, our ICONS outperforms all baselines across different selection ratios, and remarkably achieves above-full-dataset performance, surpassing the original datasets at a 60% selection ratio for LLaVA-665K. Importantly, the selected subset shows strong transferability, e.g., LLaVA-ICONS-133K maintains 95.5-113.9% relative performance across unseen tasks, suggesting that ICONS identifies fundamentally valuable training data. We summarize our key contributions:
@@
    We propose ICONS, a simple yet effective method for multitask vision-language data selection that identifies broadly valuable training examples via majority voting over task-specific gradient influence scores.
@@
    Our consensus-based selection outperforms all baselines (§3.2) and we ablate influence aggregation strategies and show the advantage of voting-based consensus (§3.3). We further show that ICONS exceeds 102% of full-dataset performance at a 60% selection ratio on LLaVA-665K3.5).
@@
    We release LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, compact 20% subsets of LLaVA-665K and Cambrian-7M respectively, achieving near-full performance (98.6% and 98.8%), transferring well to unseen tasks (§3.4), and serving as standardized training sets for resource-efficient development.
@@ 2 Influence consensus for vision-language data selection
We propose a consensus-driven, gradient-based data selection framework (Fig. 2) for visual instruction tuning datasets. We formalize the problem setup in §2.1 and establish gradient-based influence estimation preliminaries in §2.2. Our two-stage data selection framework consists of: the specialist stage (§2.3), which computes task-specific influence scores, and the generalist stage (§2.4), which builds cross-task consensus through voting-based aggregation.
@@ 2.1 Problem formulation
Given a large-scale visual instruction tuning dataset $\mathcal{D}=\{({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ containing $N$ samples, where each data point ${\bm{z}}_{i}=({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})$ includes an image ${\bm{I}}_{i}$, natural language instruction ${\bm{x}}_{i}$, and corresponding target response ${\bm{y}}_{i}$ [^1] The framework supports multi-turn conversational data, yet we formalize the problem setup for single-turn instruction-tuning for clarity and simplicity., and given access to validation data $\mathcal{V}_{k}$ for each downstream task $T_{k}\in\mathcal{T}=\{T_{1},...,T_{K}\}$, our goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes model performance across multiple downstream tasks:
$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k}),\quad\text{Rel}(f_{\mathcal{S}},T_{k})=\frac{\text{Score}(f_{\mathcal{S}},T_{k})}{\text{Score}(f_{\mathcal{D}},T_{k})},$ (1)
where $f_{\mathcal{S}}$ and $f_{\mathcal{D}}$ denote models trained on subset $\mathcal{S}$ and full dataset $\mathcal{D}$, respectively. $\text{Score}(f,T_{k})$ is the task-specific evaluation score achieved by model $f$ on task $T_{k}$. We define the average relative performance across all tasks as Rel. = $\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k})/K$. Rel. quantifies the subset-trained model’s performance relative to that of the model trained on the entire dataset, with values close to 1 indicating that the subset maintains the performance of full training coincide . Our objective is to select a subset where Rel. $\approx 1$, i.e., the model trained on the subset achieves comparable performance to the one trained with full dataset.
@@ 2.2 Preliminaries
Building on the problem formulation in §2.1, we formalize how to estimate the influence of training samples on downstream task performance. Since our goal is to maximize $\text{Rel}(f_{\mathcal{S}},T_{k})$ across tasks as defined in Eqn. 1, we need an efficient way to estimate how each training sample contributes to the $\text{Score}(f_{\mathcal{S}},T_{k})$ term in the numerator. Denote a training data point as ${\bm{z}}$ and a validation data point as ${\bm{z}}^{\prime}$ from validation set $\mathcal{V}_{k}$ for task $T_{k}$. Following pruthi2020estimating ; xia2024less , we estimate how ${\bm{z}}\in\mathcal{D}$ affects validation loss by measuring its gradient alignment with reducing validation loss on $\mathcal{V}_{k}$, which directly impacts task-specific evaluation. When training with SGD and batch size 1, using data point ${\bm{z}}$ at timestep $t$ leads to a model update ${\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}\nabla\ell({\bm{z}};{\bm{\theta}}_{t})$, where $\eta_{t}$ is the learning rate. To reduce the computational cost, we use the first-order Taylor expansion to estimate the loss on a given validation data point ${\bm{z}}^{\prime}$ at time step $t+1$ as:
@@
$\mathcal{I}_{t}({\bm{z}}\rightarrow{\bm{z}}^{\prime})$ $=\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t+1})-\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t})\approx-\eta_{t}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
which we refer to as an influence score.
The gradient-based selection approach selects training samples $\{{\bm{z}}\}$ that maximize the gradient inner product $\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle$ [^2] In practice, we use cosine similarity instead of direct inner products to avoid biasing selection toward shorter sequences, since gradient norms tend to be inversely correlated with sequence length as noted in xia2024less . through a greedy, first-order approximation, which leads to larger reductions in validation loss for point ${\bm{z}}^{\prime}$. While it omits second-order terms compared to influence functions koh2017understanding , it provides an efficient approximation to rank the impact of training samples xia2024less .
@@ 2.3 Specialist: individual task influence ranking
To rank the influence of training data for each target task, we compute the influence score of each training data point on a validation set that represents the target task distribution. Following LESS xia2024less , the process involves three steps: (1) training the model on 5% randomly selected data as a lightweight warm-up to initialize visual instruction-following capabilities, (2) computing gradients for training and validation data and compressing the gradients via random projection, and (3) computing the influence score to quantify the impact of each training data on validation set.
Step 1: Warm-up Training.  Following LESS xia2024less , we first apply LoRA hu2021lora on a small random subset $\mathcal{D}_{\mathrm{warmup}}\subset\mathcal{D}$ (5%) to obtain $f{\text{warmup}}=\text{LoRA}(f_{\text{base}},\mathcal{D}_{\text{warmup}})$. This allows the model to develop basic visual instruction-following capabilities.
@@
where $y_{i}$ and $y^{\prime}_{j}$ are the targets for $z_{i}$ and $z^{\prime}_{j}$, respectively. In order to reduce computational and storage overhead, we apply random projection to the gradient feature: $\tilde{g}_{i}=Rg_{i}$ and $\tilde{g}^{\prime}_{j}=Rg^{\prime}_{j}$, where $R\in\mathbb{R}^{d^{\prime}\times d}$ is a random projection matrix with $d^{\prime}\ll d$ that preserves inner products with high probability (johnson1984extensions, ). We further normalize the projected gradients, $\tilde{g}_{i}=\frac{\tilde{g}_{i}}{\|\tilde{g}_{i}\|_{2}},\tilde{g}^{\prime}_{j}=\frac{\tilde{g}^{\prime}_{j}}{\|\tilde{g}^{\prime}_{j}\|_{2}}$ to prevent bias from sequence length differences xia2024less .
Step 3: Influence matrix computation.  We compute the influence matrix $I\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}^{k}|}$ where each entry $I_{ij}=\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$ is the influence of the training data ${\bm{z}}_{i}$ on the validation data ${\bm{z}}^{\prime}_{j}$, and then the average influence of training data ${\bm{z}}_{i}$ on the target task $k$ is calculated as:
@@
This influence estimation process provides a task-specific ranking for the training set $\mathcal{D}$ with respect to task $\mathcal{T}_{k}$, where a higher influence score $\bar{I}_{i}$ suggests a higher influence $\mathcal{T}_{k}$.
We can select a small training subset $\mathcal{S}_{k}$ for a given task $k$ by selecting the training examples ${\bm{z}}_{i}$ with the highest-influence values $\bar{I}_{k}({\bm{z}}_{i})$. This simple greedy approach has been shown by LESS to be successful, and thus we use it as our task-specific (“specialist”) baseline. However, recall that our goal is to select a single compact subset that maximizes the performance across all tasks. We address this disconnection between task-specific rankings and overall optimization by proposing a voting-based generalist approach to identify the most broadly impactful training data.